Parth (parthkohli):

Is this problem "Lagrangable"? What is the minimum value of \(x^2 + y^2 + z^2\) given\[\dfrac{x^2}{4} + \dfrac{y^2}{5} + \dfrac{z^2}{25}=1\]and \(z=x+y\).

Parth (parthkohli):

Or just maybe...\[\dfrac{x^2}{4} + \dfrac{y^2}{5} + \dfrac{(x+y)^2}{25}=1\]

ganeshie8 (ganeshie8):

yes looks like nice candidate for lagrange :)

Parth (parthkohli):

But I'd like a solution through Lagrange Multipliers (if one exists).

Parth (parthkohli):

OK, I'm just learning about them.\[x^2 + y^2 + z^2 +\lambda (x+y-z)\]Or will we use the second condition?

Parth (parthkohli):

\[\dfrac{\partial f}{\partial x} =2x + \lambda \]Is this right?

Parth (parthkohli):

Similarly,\[\dfrac{\partial f}{\partial y}=2y + \lambda\]\[\dfrac{\partial f}{\partial z}=2z-\lambda\]

Parth (parthkohli):

OK, so \(y =z\)

ganeshie8 (ganeshie8):

hey constraint is this : \(\dfrac{x^2}{4} + \dfrac{y^2}{5} + \dfrac{(x+y)^2}{25}=1\) right ?

Parth (parthkohli):

Oh, right...

Parth (parthkohli):

So will we just remove \(z\) from existence and plug in \(x+y\) in everything?

ganeshie8 (ganeshie8):

you're given two constraints, and one function to optimize

ganeshie8 (ganeshie8):

So i think we need to merge the constraint and take it for lagrange

Parth (parthkohli):

Anyway\[x^2+y^2 + z^2+\lambda\left(\dfrac{x^2}{4} + \dfrac{y^2}{5} + \dfrac{x^2 + 2xy + y^2}{25}\right)\]

Parth (parthkohli):

Oh, forgot the \(-1\) there

ganeshie8 (ganeshie8):

^^

Parth (parthkohli):

\[x^2+y^2 + z^2+\lambda\left(\dfrac{x^2}{4} + \dfrac{y^2}{5} + \dfrac{x^2 + 2xy + y^2}{25}-1\right)\]OK.

ganeshie8 (ganeshie8):

it shouldnt matter as it becomes 0 when u take derivative..

Parth (parthkohli):

Oh whoops, didn't plug in \(x+y \) for \(z\) either.

Parth (parthkohli):

\[x^2+y^2 + (x+y)^2+\lambda\left(\dfrac{x^2}{4} + \dfrac{y^2}{5} + \dfrac{x^2 + 2xy + y^2}{25}-1\right)\]Finally

ganeshie8 (ganeshie8):

Oh, you want to eliminate z and do it in two variables is it ?

Parth (parthkohli):

Yes. :)

ganeshie8 (ganeshie8):

earlier one is also fine, but this gives less equations to solve... :)

Parth (parthkohli):

Yes, OK - should I first combine the terms or go for it?

ganeshie8 (ganeshie8):

go for it may be...

Parth (parthkohli):

Either should work. Bleh.\[\dfrac{\partial{f}}{\partial x}=2x +2x + 2y + \dfrac{\lambda}{2} + \dfrac{2\lambda}{25} + \dfrac{2\lambda y}{25}\]Phew.

ganeshie8 (ganeshie8):

looks good !

Parth (parthkohli):

Nononono not that one... should be \(2\lambda x/4\)

Parth (parthkohli):

I've confused myself. ?_?

Parth (parthkohli):

Actually, both of them. Dagnabbit.

ganeshie8 (ganeshie8):

\(\large f(x,y) = x^2+y^2 + (x+y)^2+\lambda\left(\dfrac{x^2}{4} + \dfrac{y^2}{5} + \dfrac{x^2 + 2xy + y^2}{25}-1\right) \) \(\large \dfrac{\partial{f}}{\partial x}=2x +2(x + y) + \lambda\left( \frac{2x}{4}+ \frac{2x + 2y}{25}\right)\)

Parth (parthkohli):

And\[\large \dfrac{\partial f}{\partial y}=2y + 2(x+y) +\lambda \left(\dfrac{2y}{5} + \dfrac{2y + 2x}{25}\right)\]Lesson of the day: use brackets wisely.

ganeshie8 (ganeshie8):

\(\large f(x,y) = x^2+y^2 + (x+y)^2+\lambda\left(\dfrac{x^2}{4} + \dfrac{y^2}{5} + \dfrac{x^2 + 2xy + y^2}{25}-1\right) \) \(\large \dfrac{\partial{f}}{\partial x}=2x +2(x + y) + \lambda\left( \frac{2x}{4}+ \frac{2x + 2y}{25}\right) \) \(\large \dfrac{\partial{f}}{\partial y}=2y +2(x + y) + \lambda\left( \frac{2y}{5}+ \frac{2x + 2y}{25}\right) \)

Parth (parthkohli):

And I was sitting there expanding every little thing. -_-

ganeshie8 (ganeshie8):

hahah more variables more fun :)

ganeshie8 (ganeshie8):

hey... whats the difference between \dfrac, and \frac ?

Parth (parthkohli):

`\dfrac` doesn't reduce the numbers in fractions to half their original size.

ganeshie8 (ganeshie8):

oh ! \(\large \frac{1}{2}, \dfrac{1}{2}\) gotcha :)

Parth (parthkohli):

Like you can see how the fractions come out really small in `\frac`. By the way, fractions shouldn't be coming out small if you write \(\LaTeX\) within `\[ \]`. I see that you use `\( \)`

Parth (parthkohli):

All right, let's subtract the two equations.

Parth (parthkohli):

\[\dfrac{2x}{4} - \dfrac{2y}{5}=0\]

Parth (parthkohli):

Oh wait.

Parth (parthkohli):

\[\left( 2x -2y \right) +\left( \dfrac{x}2 - \dfrac{2y}{5}\right)=0\]

Parth (parthkohli):

\(\LaTeX\), Y U NO WORK

Parth (parthkohli):

Oh, looks like there should be a \(\lambda\) in there too?

ganeshie8 (ganeshie8):

yes lol

Parth (parthkohli):

\[\left( 2x -2y \right) +\lambda \left( \dfrac{x}2 - \dfrac{2y}{5}\right)=0\]

ganeshie8 (ganeshie8):

the better way is not to subtract the equations

Parth (parthkohli):

How do we get rid of the \(\lambda\)? ;_;

ganeshie8 (ganeshie8):

cuz, by subtracting the equations we lost some info

ganeshie8 (ganeshie8):

Lagrange works like this : At local max/min points, the gradiants are proportional, so we get : 1) \(\dfrac{\partial f}{\partial x} = 0\) 2) \(\dfrac{\partial f}{\partial y} = 0\)

ganeshie8 (ganeshie8):

above both give u two equations in 3 variables : \(\large x, y, \lambda\)

ganeshie8 (ganeshie8):

but how can u solve 3 variables wid just 2 equations ?

Parth (parthkohli):

Oh, so we should have included the \(z\) :O

ganeshie8 (ganeshie8):

use the constraint as 3rd equation : 3) \(\dfrac{x^2}{4} + \dfrac{y^2}{5} + \dfrac{(x+y)^2}{25}=1\)

Parth (parthkohli):

Oh, OK.

ganeshie8 (ganeshie8):

3 equations and 3 unknowns, feed it to wolfram. dont bother solving it manually lol

ganeshie8 (ganeshie8):

it gets messy

Parth (parthkohli):

OK, thank you for helping! ^_^ The way I learnt the basics was this: https://www.quora.com/What-is-a-Lagrange-multiplier

Parth (parthkohli):

Can you explain the "gradients are proportional" part?

ganeshie8 (ganeshie8):

once u have 3 equations, its just algebra to sovle them..

ganeshie8 (ganeshie8):

yeah sure :)

Parth (parthkohli):

Aah, I'm beginning to see why \(\partial/\partial x\) is 0.

Parth (parthkohli):

It's just like how any one-variable function works.

ganeshie8 (ganeshie8):

two different things : 1) finding max/min of a multivariable funciton : f(x, y, z) 2) finding max/min of a multivariable funciton : f(x, y, z) under given constraints

Parth (parthkohli):